40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Dec 1 17:16:31 2015, gautam, Update, CDS, Script for copying BLRMS filters makeFilterFiles.bash.zip
    Reply  Wed Dec 2 18:54:20 2015, gautam, Update, CDS, Changes to C1MCS, C1PEM  
       Reply  Thu Dec 3 18:18:48 2015, gautam, Update, CDS, BLRMS for optics suspensions - library block 
          Reply  Fri Dec 4 18:20:36 2015, gautam, Update, CDS, BLRMS for IMC setup c1pem_timing.pngC1SUS_overclocked.png
             Reply  Mon Dec 7 00:45:28 2015, ericq, Update, CDS, daqd is mad 
                Reply  Mon Dec 7 10:45:21 2015, ericq, Update, CDS, daqd is mad 
                   Reply  Mon Dec 7 11:25:10 2015, jamie, Update, CDS, daqd is mad 
                   Reply  Mon Dec 14 23:56:29 2015, ericq, Update, CDS, c1pem reverted 
          Reply  Tue Apr 5 14:16:34 2016, gautam, Update, CDS, BLRMS for optics suspensions - library block UPDATED 
             Reply  Wed Nov 2 22:56:45 2016, gautam, Update, CDS, c1pem revamped 
                Reply  Thu Nov 3 12:38:42 2016, gautam, Update, CDS, c1pem revamped BLRMSresp.pdf
                   Reply  Thu Nov 3 13:36:16 2016, ericq, Update, CDS, c1pem revamped 
                   Reply  Fri Jan 27 21:21:35 2017, gautam, Update, CDS, c1pem revamped 
Message ID: 11844     Entry time: Thu Dec 3 18:18:48 2015     In reply to: 11840     Reply to this: 11849   12064
Author: gautam 
Type: Update 
Category: CDS 
Subject: BLRMS for optics suspensions - library block 

In order to be consistent with the naming conventions for the new BLRMS filters, I made a library block that takes all the input signals of interest (i.e. for a generic optic, the coil signals, the local damping shadow sensors, and the Oplev Pitch and Yaw signals - so a total of 12 signals, unused ones can be grounded). The block is called "sus_single_BLRMS". Inside the block, I've put in 12 BLRMS library blocks, with each input signal going to one of them. All the 7 outputs of the BLRMS block are terminated (I got a compiling error if I did not do this). The idea is to identify the optic using this block, e.g. MC2_BLRMS. The BLRMS filters inside are called UL_COIL, UR_COIL etc, so the BLRMS channels will end up being called C1:SUS-MC2_BLRMS_UL_COIL_0p01_0p03 and so on. I tried implementing this in C1PEM, but immediately after compiling and restarting the model, I noticed some strange behaviour in the seismic rainbow STS strip in the control room - this was right after the model was restarted, before I attempted to make any changes to the C1PEM.txt file and add filters. I then manually opened up the filter bank screens for the RMS_STS1Z bandpass and lowpass filters, and saw that the filter switches were OFF - I wonder if this has something to do with these settings not being updated in the SDF tables? So I manually turned them on and cleared the filter hitsory for all 7 low pass and band pass filter banks, but the traces on the seismic striptool did not return to their nominal levels. I manually checked the filter shapes with Foton and they seem alright. Anyways, for now, I've reverted to the C1PEM model before I made any changes, and the seismic strip looks to be back at its normal level - when I recompiled and restarted the model with the changes I made removed, the STS1Z BLRMS bandpass and lowpass filters were ON by default again! I'm not sure what I'm doing wrong here, I will investigate this further. 

ELOG V3.1.3-